# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v2

@@ 3.2 Two-Stage PG-DPO Algorithm
11:     Simulate path with $(\pi_{\theta^{*}},C_{\phi^{*}})$ using $N^{\prime}$ steps
@@
13:     $\qquad\displaystyle J^{(j)}\leftarrow\sum_{n=0}^{N^{\prime}-1}e^{-\delta t_{n}}U\bigl{(}C_{t_{n}}^{(j)}\bigr{)}\,\Delta t^{\prime}+\kappa e^{-\delta T}U\bigl{(}X_{T}^{(j)}\bigr{)}$
14:     Evaluate $\lambda_{t}^{(j)}$, $\partial_{x}\lambda_{t}^{(j)}$, $\partial_{\bm{Y}}\lambda_{t}^{(j)}$ and append
@@
16:Compute $\hat{\lambda}_{t},\widehat{\partial_{x}\lambda}_{t},\widehat{\partial_{\bm{Y}}\lambda}_{t}$triangleright E.g., by averaging estimates from loop
@@
This Two-Stage approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly from the PMP first-order conditions, bypassing the need for fully converged policy networks for deployment. The procedure consists of two distinct stages, described below using a single level of enumeration:
@@
    After completing the $K_{0}$ warm-up iterations, the trained policy networks $(\bm{\pi}_{\theta^{*}},C_{\phi^{*}})$ are typically not directly used for generating controls during deployment. Instead, to determine the control actions for any given state $(t_{k},X_{k},\mathbf{Y}_{k})$ encountered during deployment, we employ a Monte Carlo-like averaging procedure to obtain reliable costate estimates needed for the PMP formulas:
    First, to obtain the stabilized costate estimates corresponding to the state $(t_{k},X_{k},\mathbf{Y}_{k})$, we perform the following averaging: Simulate multiple ($M_{\text{MC}}$) forward paths starting from $(t_{k},X_{k},\mathbf{Y}_{k})$ (or states very close to it) up to the horizon $T$, using the fixed policy parameters $(\theta^{*},\phi^{*})$ obtained from Stage 1. For each simulated path $j=1,\dots,M_{\text{MC}}$, calculate its realized reward $J^{(j)}(t_{k},X_{k},\mathbf{Y}_{k};\theta^{*},\phi^{*})$ and apply BPTT to obtain the costate estimates $\lambda_{k}^{(j)},\partial_{x}\lambda_{k}^{(j)},\partial_{\mathbf{Y}}\lambda_{k}^{(j)}$ specifically at the starting node $k$ for that path $j$. The final “stabilized” estimates are then computed by averaging these values across the $M_{\text{MC}}$ paths:
    $\lambda_{k}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\lambda_{k}^{(j)},\quad\partial_{x}\lambda_{k}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{x}\lambda_{k}^{(j)},\quad\partial_{\mathbf{Y}}\lambda_{k}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{\mathbf{Y}}\lambda_{k}^{(j)}$
@@
    Next, we plug these averaged costate estimates ($\lambda_{k},\partial_{x}\lambda_{k},\partial_{\mathbf{Y}}\lambda_{k}$) directly into the analytical formulas derived from the Pontryagin Maximum Principle’s first-order conditions (Section 2.2). The optimal consumption is calculated as:
    $C_{k}^{\mathrm{PMP}}=\bigl{(}U^{\prime}\bigr{)}^{-1}\!\bigl{(}e^{\delta\,t_{k}}\,\lambda_{k}\bigr{)}.$
    The optimal investment portfolio is calculated as:
    $\bm{\pi}_{k}^{\mathrm{PMP}}$ $=-\frac{1}{X_{k}\,\partial_{x}\lambda_{k}}\,\Sigma^{-1}\!\bigl{(}t_{k},\mathbf{Y}_{k}\bigr{)}\Bigl{\{}\lambda_{k}\bigl{[}\bm{\mu}(t_{k},\mathbf{Y}_{k})-r(t_{k},\mathbf{Y}_{k})\mathbf{1}\bigr{]}$ (14)
    $\qquad+\bm{\sigma}(t_{k},\mathbf{Y}_{k})^{\top}(\partial_{\mathbf{Y}}\lambda_{k})\,\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\rho^{\top}\Psi^{-1}\Bigr{\}}.$ (15)
    Finally, these analytically computed values $(C_{k}^{\mathrm{PMP}},\bm{\pi}_{k}^{\mathrm{PMP}})$, based on the averaged costate estimates, are used as the control actions for the state $(t_{k},X_{k},\mathbf{Y}_{k})$ during deployment.
In summary, the Two-Stage PG-DPO approach performs a short warm-up training (Stage 1) using the baseline PG-DPO process to obtain a policy $(\theta^{*},\phi^{*})$ whose BPTT generates reasonably stable costate estimates. It then leverages these estimates in Stage 2, refining them through Monte Carlo averaging at the point of deployment and plugging them into the known analytical structure of the optimal control derived from PMP. This strategy can significantly reduce the overall computational burden while yielding high-quality, near-optimal controls. The Monte Carlo averaging step in Stage 2 is crucial for obtaining reliable inputs for the PMP formulas from the potentially noisy single-path estimates.
3.3 Rationale for the Two-Stage Approach
3.3.1 The Parabolic Nature of Financial Models and PDE Regularity
@@
(U) Uniformly Elliptic Uniformly positive definite across the state space The diffusion coefficients $\bm{\sigma}$ (for wealth $X_{t}$) and $\bm{\sigma}_{Y}$ (for state $\mathbf{Y}_{t}$) are such that the resulting diffusion matrix in the HJB is non-zero and bounded below by a positive constant everywhere ($\bm{\sigma}\Psi\bm{\sigma}^{\top}$ and $\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}$ uniformly positive definite). Drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) are also sufficiently regular (e.g., Lipschitz).
@@
(H) Hypo-elliptic / Uniform on Compacts May degenerate at some points, but regularity is achieved via system structure (e.g., drift-diffusion coupling via Hörmander’s condition) or holds within a compact domain of interest. Degeneracy occurs away from boundaries. Diffusion coefficients ($\bm{\sigma}$ or $\bm{\sigma}_{Y}$) can become zero at certain points within the domain. However, interaction with non-zero drift terms or other diffusion terms maintains regularity.
@@
In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are modeled by stochastic differential equations (SDEs), the Value Function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically solves a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming & Soner, , 2006; Yong & Zhou, , 2012). This HJB equation exhibits a parabolic form. The specific nature of its parabolicity is primarily determined by the diffusion coefficients of the underlying SDEs, namely $\bm{\sigma}$ (related to assets) and $\bm{\sigma}_{Y}$ (related to state variables). While the drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) influence the first-order terms in the HJB and are crucial for the solution’s properties, they do not dictate the type’ of parabolicity, which is mainly governed by the second-order terms (diffusion).
The behavior of the diffusion matrix $a(t,x,\mathbf{y})$ (the matrix multiplying second-order derivatives in the HJB equation) allows for classifying parabolic PDEs into three main classes (see, e.g., Evans, , 2022). This classification significantly impacts the regularity (smoothness) of the Value Function $V$ and its associated Costates (the first-order derivatives $\nabla V$) (cf. Evans, , 2022, Ch. 6-7). Costates represent the sensitivity of the Value Function, and Pontryagin’s Maximum Principle links optimal Costates to the optimal policy. The core idea of Two-Stage PG-DPO involves estimating Costates and their derivatives via BPTT and using these estimates in the PMP formulas. Therefore, the regularity of the Costates—tied to the model’s parabolicity class—directly affects the difficulty of Costate estimation and the effectiveness of the Two-Stage method.
The table 1 summarizes these classes, their properties, links to SDE coefficients, and relevant examples within the dynamic portfolio optimization context.

### v3

@@ 3.2 Projected PG-DPO (P-PGDPO) Algorithm
11:     Simulate path from $(t,X,\bm{Y})$ with $(\pi_{\theta^{*}},C_{\phi^{*}})$ using $N^{\prime}$ steps. Let $t_{0}^{\prime}=t$.
@@
13:     $\qquad\displaystyle J^{(j)}\leftarrow\sum_{n=0}^{N^{\prime}-1}e^{-\delta(t_{n}-t_{0}^{\prime})}U\bigl{(}C_{t_{n}}^{(j)}\bigr{)}\,\Delta t^{\prime}+\kappa e^{-\delta(T-t_{0}^{\prime})}U\bigl{(}X_{T}^{(j)}\bigr{)}$
14:     Evaluate $\lambda_{t}^{(j)}$, $\partial_{x}\lambda_{t}^{(j)}$, $\partial_{\bm{Y}}\lambda_{t}^{(j)}$ (costates at $t_{0}^{\prime}=t$) via BPTT on $J^{(j)}$ and append to lists $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}$.
@@
16:Compute $\hat{\lambda}_{t},\widehat{\partial_{x}\lambda}_{t},\widehat{\partial_{\bm{Y}}\lambda}_{t}$triangleright E.g., by averaging estimates from lists $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}$
@@
This P-PGDPO approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly from the PMP first-order conditions, bypassing the need for fully converged policy networks for deployment. The procedure consists of two distinct stages, described below using a single level of enumeration:
@@
    After completing the $K_{0}$ warm-up iterations, the trained policy networks $(\bm{\pi}_{\theta^{*}},C_{\phi^{*}})$ are typically not directly used for generating controls during deployment. Instead, to determine the control actions for any given state $(t_{k},X_{k},\mathbf{Y}_{k})$ encountered during deployment (let this deployment time be $t_{k}=t_{deploy}$), we employ a Monte Carlo-like averaging procedure to obtain reliable costate estimates needed for the PMP formulas:
    First, to obtain the stabilized costate estimates corresponding to the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$, we perform the following averaging: Simulate multiple ($M_{\text{MC}}$) forward paths starting from $(t_{deploy},X_{k},\mathbf{Y}_{k})$ (or states very close to it) up to the horizon $T$, using the fixed policy parameters $(\theta^{*},\phi^{*})$ obtained from Stage 1. For each simulated path $j=1,\dots,M_{\text{MC}}$, calculate its realized reward $J^{(j)}(t_{deploy},X_{k},\mathbf{Y}_{k};\theta^{*},\phi^{*})$ (using discounting relative to $t_{deploy}$ as in Algorithm 2) and apply BPTT to obtain the costate estimates $\lambda_{deploy}^{(j)},\partial_{x}\lambda_{deploy}^{(j)},\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$ specifically at the starting node $t_{deploy}$ for that path $j$. The final “stabilized” estimates are then computed by averaging these values across the $M_{\text{MC}}$ paths:
    $\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\lambda_{deploy}^{(j)},\quad\partial_{x}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{x}\lambda_{deploy}^{(j)},\quad\partial_{\mathbf{Y}}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$
@@
    Next, we plug these averaged costate estimates ($\lambda_{deploy},\partial_{x}\lambda_{deploy},\partial_{\mathbf{Y}}\lambda_{deploy}$) directly into the analytical formulas derived from the Pontryagin Maximum Principle’s first-order conditions (Section 2.2). The optimal consumption at $t_{deploy}$ is calculated as:
    $C_{deploy}^{\mathrm{PMP}}=\bigl{(}U^{\prime}\bigr{)}^{-1}\!\bigl{(}e^{\delta\,t_{deploy}}\,\lambda_{deploy}\bigr{)}.$
    The optimal investment portfolio at $t_{deploy}$ is calculated as:
    $\bm{\pi}_{deploy}^{\mathrm{PMP}}$ $=-\frac{1}{X_{k}\,\partial_{x}\lambda_{deploy}}\,\Sigma^{-1}\!\bigl{(}t_{deploy},\mathbf{Y}_{k}\bigr{)}\Biggl{\{}\lambda_{deploy}\bigl{[}\bm{\mu}(t_{deploy},\mathbf{Y}_{k})-r(t_{deploy},\mathbf{Y}_{k})\mathbf{1}\bigr{]}$
    $\qquad+\bigl{(}\bm{\sigma}(t_{deploy},\mathbf{Y}_{k})\,\bm{\rho}\,\bm{\sigma}_{Y}(t_{deploy},\mathbf{Y}_{k})\bigr{)}(\partial_{\mathbf{Y}}\lambda_{deploy})\Biggr{\}}.$ (15)
    (This formula aligns with Eq. (11), where $\lambda_{deploy}$ corresponds to $V_{x}$, $\partial_{x}\lambda_{deploy}$ to $V_{xx}$, and $\partial_{\mathbf{Y}}\lambda_{deploy}$ to $V_{x\mathbf{Y}}$ at state $(t_{deploy},X_{k},\mathbf{Y}_{k})$.)
    Finally, these analytically computed values $(C_{deploy}^{\mathrm{PMP}},\bm{\pi}_{deploy}^{\mathrm{PMP}})$, based on the averaged costate estimates, are used as the control actions for the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$ during deployment.
In summary, the Projected PG-DPO (P-PGDPO) approach performs a short warm-up training (Stage 1) using the baseline PG-DPO process to obtain a policy $(\theta^{*},\phi^{*})$ whose BPTT generates reasonably stable costate estimates. It then leverages these estimates in Stage 2, refining them through Monte Carlo averaging at the point of deployment and plugging them into the known analytical structure of the optimal control derived from PMP. This strategy can significantly reduce the overall computational burden while yielding high-quality, near-optimal controls. The Monte Carlo averaging step in Stage 2 is crucial for obtaining reliable inputs for the PMP formulas from the potentially noisy single-path estimates.
3.3 Rationale for the Projected PG-DPO (P-PGDPO) Approach
In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are modeled by stochastic differential equations (SDEs), the Value Function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically solves a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming and Soner, , 2006; Yong and Zhou, , 2012). This HJB equation exhibits a parabolic form. The specific nature of its parabolicity is primarily determined by the diffusion coefficients of the underlying SDEs, namely $\bm{\sigma}$ (related to assets) and $\bm{\sigma}_{Y}$ (related to state variables). While the first-order drift terms (involving $\bm{\mu},\bm{\mu}_{Y}$) are crucial for the solution’s properties, they do not affect the parabolicity classification, which is mainly governed by the second-order diffusion terms.
The behavior of the diffusion matrix $a(t,x,\mathbf{y})$ (the matrix multiplying second-order derivatives in the HJB equation) allows for classifying parabolic PDEs into three main classes (see, e.g., Evans, , 2022). This classification significantly impacts the regularity (smoothness) of the Value Function $V$ and its associated Costates (the first-order derivatives $\nabla V$) (cf. Evans, , 2022, Ch. 6-7). Costates represent the sensitivity of the Value Function, and Pontryagin’s Maximum Principle links optimal Costates to the optimal policy. The core idea of Projected PG-DPO (P-PGDPO) involves estimating Costates and their derivatives via BPTT and using these estimates in the PMP formulas. Therefore, the regularity of the Costates—tied to the model’s parabolicity class—directly affects the difficulty of Costate estimation and the effectiveness of the P-PGDPO method.
@@ 3.3 Rationale for the Projected PG-DPO (P-PGDPO) Approach
(U) Uniformly Elliptic Uniformly positive definite across the state space The diffusion coefficients $\bm{\sigma}$ (for wealth $X_{t}$) and $\bm{\sigma}_{Y}$ (for state $\mathbf{Y}_{t}$) are such that the resulting diffusion matrix in the HJB is non-zero and bounded below by a positive constant everywhere (e.g., $\bm{\sigma}\Psi\bm{\sigma}^{\top}$ and $\operatorname{diag}(\bm{\sigma}_{Y})\Phi\operatorname{diag}(\bm{\sigma}_{Y})^{\top}$ uniformly positive definite). Drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) are also sufficiently regular (e.g., Lipschitz).
@@
(H) Hypo-elliptic May degenerate (even globally). Regularity is achieved if the system structure satisfies conditions like Hörmander’s (i.e., Lie brackets of vector fields span the tangent space). The problem might be uniformly elliptic on compact subsets of interest. Diffusion coefficients ($\bm{\sigma}$ or $\bm{\sigma}_{Y}$) can become zero at certain points within the domain. However, interaction with non-zero drift terms or other diffusion terms maintains regularity.